OpenStudy (anonymous):

Finding the solutions f(x)=x^2+6x+1

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

(x-3)^2 - (√8)^2= (x-3-√8)(x-3+√8)= (x-3-2√2)(x-3+2√2)

OpenStudy (anonymous):

i got u qurl <3

OpenStudy (anonymous):

\[f(x)=x^2+6x+1 \] solutions means i assume solve \[x^2+6x+1=0\] right?

OpenStudy (anonymous):

you and either use the quadratic formula or complete the square second method is probably easiest here because of the \(6x\)

OpenStudy (anonymous):

\[x^2+6x=1\\ x^2+6x=-1\\ (x+3)^2=-1+9=8\\ x+3-\pm\sqrt8\\ x=-3\pm\sqrt8\]

OpenStudy (anonymous):

or even \[x=3\pm2\sqrt2\]